### v1

@@ 4.6 Optimal Treatment Assignment with Interval-Identified ATE
$b_{L}(p)\equiv\max_{1\leq k\leq K_{L}}\{\tilde{\mathfrak{l}}_{k}+\mathfrak{l}_{k}p\}\leq ATE\leq\min_{1\leq k\leq K_{U}}\{\tilde{\mathfrak{u}}_{k}+\mathfrak{u}_{k}p\}\equiv b_{U}(p)$
for some fixed and known $K_{L}$, $K_{U}$, $\tilde{\mathfrak{l}}_{1},\ldots,\tilde{\mathfrak{l}}_{K_{L}}$, $\tilde{\mathfrak{u}}_{1},\ldots,\tilde{\mathfrak{u}}_{K_{U}}$, $\mathfrak{l}_{1},\ldots,\mathfrak{l}_{K_{L}}$ and $\mathfrak{u}_{1},\ldots,\mathfrak{u}_{K_{U}}$. [^7] Although Christensen et al., (2023) do not write the form of their bounds as they are written here, the representation here is equivalent to the one in that paper upon proper definition of $p$ since the elements in the intersection bounds are smooth functions of a reduced-form parameter.
@@
Suppose $\widehat{\mathcal{D}}=\{\hat{d}\}$, where $\hat{d}$ is defined by (4.12). Then Assumptions 3.2 and 3.6 are satisfied for
$\hat{\gamma}_{L}(d)=\hat{\gamma}_{U}(d)=(\varepsilon_{1}^{\prime},\ldots,\varepsilon_{m}^{\prime},\underline{k}_{1},\ldots,\underline{k}_{m},\bar{k}_{1},\ldots,\bar{k}_{m},s_{1}^{\mathfrak{l}},\ldots,s_{m}^{\mathfrak{l}},s_{1}^{\mathfrak{u}},\ldots,s_{m}^{\mathfrak{u}})^{\prime},$
where $\underline{k}_{i}=\operatorname{argmin}_{1\leq k\leq K_{U}}\{\tilde{\mathfrak{u}}_{k}+\mathfrak{u}_{k}(\hat{p}+\varepsilon_{i})\}$, $\bar{k}_{i}=\operatorname*{argmax\;}_{1\leq k\leq K_{U}}\{\tilde{\mathfrak{l}}_{k}+\mathfrak{l}_{k}(\hat{p}+\varepsilon_{i})\}$, $s_{i}^{\mathfrak{l}}=\operatorname*{sign}(\tilde{\mathfrak{l}}_{\bar{k}_{i}}+\mathfrak{l}_{\bar{k}_{i}}(\hat{p}+\varepsilon_{i}))$ and $s_{i}^{\mathfrak{u}}=\operatorname*{sign}(\tilde{\mathfrak{u}}_{\underline{k}_{i}}+\mathfrak{u}_{\underline{k}_{i}}(\hat{p}+\varepsilon_{i}))$ for $i=1,\ldots,m$.
5 Conditional Confidence Intervals
We now generalize the conditional CI construction described in Section 2.2 to apply in the general framework of Section 3, covering all example applications discussed above. As in Section 2.2, we are interested in forming probabilistic lower and upper bounds $\widehat{L}(\hat{d})_{\alpha}^{C}$ and $\widehat{U}(\hat{d})_{1-\alpha}^{C}$ that satisfy (2.2) and (2.5) for all $d\in\{0,\ldots,T\}$ as endpoints in the formation of a conditionally valid CI because the researcher’s interest in inference on option $d$ only arises when it is a member of the estimated set $\widehat{\mathcal{D}}$.
To begin, we characterize the conditional distributions of $\widehat{L}(d)$ given $d\in\widehat{\mathcal{D}}$, $\hat{j}_{L}(d)=j_{L}^{*}$ and $\hat{\gamma}_{L}(d)=\gamma_{L}^{*}$ and $\widehat{U}(d)$ given $d\in\widehat{\mathcal{D}}$, $\hat{j}_{U}(d)=j_{U}^{*}$ and $\hat{\gamma}_{U}(d)=\gamma_{U}^{*}$. These conditional distributions depend upon the nuisance parameter $p$. As a first step, form sufficient statistics for $p$ that are asymptotically independent of $\widehat{L}(d)$ given $\hat{j}_{L}(d)=j_{L}^{*}$ and $\widehat{U}(d)$ given $\hat{j}_{U}(d)=j_{U}^{*}$:
$\widehat{Z}_{L}(d,j_{L}^{*})=\sqrt{n}\hat{p}-\hat{b}_{L}(d,j_{L}^{*})\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p})$
@@ 5 Conditional Confidence Intervals
$\widehat{Z}_{U}(d,j_{U}^{*})=\sqrt{n}\hat{p}-\hat{b}_{U}(d,j_{U}^{*})\sqrt{n}(\tilde{u}_{d,j_{U}^{*}}+u_{d,j_{U}^{*}}\hat{p})$
@@
$\hat{b}_{L}(d,j_{L}^{*})=\widehat{\Sigma}\ell_{d,j_{L}^{*}}^{\prime}\left(\ell_{d,j_{L}^{*}}\widehat{\Sigma}\ell_{d,j_{L}^{*}}^{\prime}\right)^{-1}\quad\text{and}\quad\hat{b}_{U}(d,j_{U}^{*})=\widehat{\Sigma}u_{d,j_{U}^{*}}^{\prime}\left(u_{d,j_{U}^{*}}\widehat{\Sigma}u_{d,j_{U}^{*}}^{\prime}\right)^{-1}.$
Under Assumption 3.2, Lemma 1 of McCloskey, (2024) implies
@@
$\sim\left.\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p})\right|\left\{\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d,j_{L}^{*}),d,j_{L}^{*},\gamma_{L}^{*}\right)\leq\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p})\leq\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}(d,j_{L}^{*}),d,j_{L}^{*},\gamma_{L}^{*}\right),\right.$
$\left.\widehat{\mathcal{V}}_{L}^{0}\left(\widehat{Z}_{L}(d,j_{L}^{*}),d,j_{L}^{*},\gamma_{L}^{*}\right)\geq 0\right\}$
@@
$\sim\sqrt{n}(\tilde{u}_{d,j_{U}^{*}}+u_{d,j_{U}^{*}}\hat{p})\left\{\widehat{\mathcal{V}}_{U}^{-}\left(\widehat{Z}_{U}(d,j_{U}^{*}),d,j_{U}^{*},\gamma_{U}^{*}\right)\leq\sqrt{n}(\tilde{u}_{d,j_{U}^{*}}+u_{d,j_{U}^{*}}\hat{p})\leq\widehat{\mathcal{V}}_{U}^{+}\left(\widehat{Z}_{U}(d,j_{U}^{*}),d,j_{U}^{*},\gamma_{U}^{*}\right),\right.$
$\left.\widehat{\mathcal{V}}_{U}^{0}\left(\widehat{Z}_{U}(d,j_{U}^{*}),d,j_{U}^{*},\gamma_{U}^{*}\right)\geq 0\right\}$
@@
$\widehat{\mathcal{V}}_{M}^{-}(z,d,j,\gamma)=\max_{k:(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}<0}\frac{\sqrt{n}(c^{M}(d,j,\gamma))_{k}-(A^{M}(d,j,\gamma)z)_{k}}{(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}},$
$\widehat{\mathcal{V}}_{M}^{+}(z,d,j,\gamma)=\min_{k:(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}>0}\frac{\sqrt{n}(c^{M}(d,j,\gamma))_{k}-(A^{M}(d,j,\gamma)z)_{k}}{(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}},$
$\widehat{\mathcal{V}}_{M}^{0}(z,d,j,\gamma)=\min_{k:(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}=0}\sqrt{n}(c^{M}(d,j,\gamma))_{k}-(A^{M}(d,j,\gamma)z)_{k}$
@@
Under Assumptions 3.3 and 3.4, the distribution of $\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p})$ can be approximated by a $\mathcal{N}(\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}p),\ell_{d,j_{L}^{*}}\Sigma\ell_{d,j_{L}^{*}}^{\prime})$ distribution asymptotically and $\widehat{Z}_{L}(d,j_{L}^{*})$ is asymptotically independent. For $\alpha\in(0,1)$, define $\widehat{L}(d)_{\alpha}^{C}$ to solve
$F_{TN}\left(\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p});\mu,\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime},\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right),\right.$
$\left.\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{Z}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right)\right)=1-\alpha$
@@
$F_{TN}\left(\sqrt{n}(\tilde{u}_{d,\hat{j}_{U}(d)}+u_{d,\hat{j}_{U}(d)}\hat{p});\mu,u_{d,\hat{j}_{U}(d)}\widehat{\Sigma}u_{d,\hat{j}_{U}(d)}^{\prime},\widehat{\mathcal{V}}_{U}^{-}\left(\widehat{Z}_{U}(d,\hat{j}_{U}(d)),d,\hat{j}_{U}(d),\hat{\gamma}_{U}(d)\right),\right.$
$\left.\widehat{\mathcal{V}}_{U}^{+}\left(\widehat{Z}_{U}(d,\hat{j}_{U}(d)),d,\hat{j}_{U}(d),\hat{\gamma}_{U}(d)\right)\right)=1-\alpha$
in $\mu$. Results in Pfanzagl, (1994) imply that $\widehat{L}(d)_{\alpha}^{C}$ and $\widehat{U}(d)_{\alpha}^{C}$ are optimal $\alpha$ quantile-unbiased estimators of $\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}p)$ and $\sqrt{n}(\tilde{u}_{d,j_{L}^{*}}+u_{d,j_{L}^{*}}p)$ asymptotically.
@@
Suppose Assumptions 3.13.5 hold. Then, for any $d\in\{0,\ldots,T\}$ and $20<\alpha_{1},\alpha_{2}<1/$ 2,
$\liminf_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}_{n}}\left\{\left[\mathbb{P}\left(\left.[L(d),U(d)]\subseteq\left(n^{-1/2}\widehat{L}(d)_{\alpha_{1}}^{C},n^{-1/2}\widehat{U}(d)_{1-\alpha_{2}}^{C}\right)\right|d\in\widehat{\mathcal{D}}\right)-(1-\alpha_{1}-\alpha_{2})\right]\right.$
$\left.\cdot\mathbb{P}(d\in\widehat{\mathcal{D}})\right\}\geq 0$
for all $d\in\{0,\ldots,T\}$.
6 Unconditional Confidence Intervals
In parallel with the previous section, we now generalize the unconditional CI constructions described in Section 2.3 to the general framework of Section 3. Here, we would like to construct CIs that unconditionally cover the identified interval corresponding to a unique data-dependent object of inferential interest. As mentioned in Section 2.3 if only unconditional coverage of $[L(d),U(d)]$ is desired, the conditional confidence interval (5.1) can be unnecessarily wide. We describe two different methods to form the unconditional probabilistic bounds that constitute the endpoints of these unconditional CIs in this general framework.
The general formation of the probabilistic bounds based upon projecting simultaneous confidence bounds for all possible values of $L(\hat{d})$ and $U(\hat{d})$ proceeds by computing $\hat{c}_{1-\alpha,M}$, the $1-\alpha$ quantile of $\max_{i\in\{1,\ldots,(T+1)J_{M}\}}\hat{\zeta}_{M,i}/\sqrt{\widehat{\Sigma}_{M,i}}$, where $\hat{\zeta}_{M}\sim\mathcal{N}(0,\widehat{\Sigma}_{M})$ for $M=L,U$ with $\widehat{\Sigma}_{L}=\ell^{mat}\widehat{\Sigma}\ell^{mat\prime}$, $\widehat{\Sigma}_{U}=u^{mat}\widehat{\Sigma}u^{mat\prime}$, $\ell^{mat}=(\ell_{0,1},\ldots,\ell_{0,J_{L}},\ldots,\ell_{T,1},\ldots,\ell_{T,J_{L}})^{\prime}$ and $u^{mat}=(u_{0,1},\ldots,u_{0,J_{U}},\ldots,u_{T,1},\ldots,u_{T,J_{U}})^{\prime}$. The lower level $1-\alpha$ projection confidence bound for $\sqrt{n}L(\hat{d})$ is $\widehat{L}(\hat{d})_{\alpha}^{P}=\sqrt{n}\widehat{L}(\hat{d})-\sqrt{\widehat{\Sigma}_{L,\hat{d}J_{L}+\hat{j}_{L}(\hat{d})}}\hat{c}_{1-\alpha,L}$ and the upper level $1-\alpha$ projection confidence bound for $\sqrt{n}U(\hat{d})$ is $\widehat{U}_{1-\alpha}^{P}(\hat{d})=\sqrt{n}\widehat{U}(\hat{d})+\sqrt{\widehat{\Sigma}_{U,\hat{d}J_{U}+\hat{j}_{U}(\hat{d})}}\hat{c}_{1-\alpha,U}$.

### v2

@@ 4.4 Optimal Treatment Assignment with Interval-Identified ATE
$b_{L}(p)\equiv\max_{j\in\{1,\ldots,J_{L}\}}\{\tilde{\ell}_{j}+\ell_{j}p\}\leq ATE\leq\min_{j\in\{1,\ldots,J_{U}\}}\{\tilde{u}_{k}+u_{k}p\}\equiv b_{U}(p)$
for some fixed and known $J_{L}$, $J_{U}$, ~$\tilde{\ell}_{1},\ldots,\tilde{\ell}_{J_{L}}$, $\tilde{u}_{1},\ldots,\tilde{u}_{J_{U}}$, $\ell_{1},\ldots,\ell_{J_{L}}$ and $u_{1},\ldots,u_{J_{U}}$. [^7] Although Christensen et al., (2023) do not write the form of their bounds as they are written here, the representation here is equivalent to the one in that paper upon proper definition of $p$ since the elements in the intersection bounds are smooth functions of a reduced-form parameter. Therefore, Assumption 3.1 trivially holds for the ATE.
@@
Suppose $\widehat{\mathcal{D}}=\{\hat{d}\}$, where $\hat{d}$ is defined by (4.10). Then Assumptions 3.2 and 3.3 are satisfied for
$\hat{\gamma}_{L}(d)=\hat{\gamma}_{U}(d)=(\varepsilon_{1}^{\prime},\ldots,\varepsilon_{m}^{\prime},\underline{k}_{1},\ldots,\underline{k}_{m},\bar{k}_{1},\ldots,\bar{k}_{m},s_{1}^{\ell},\ldots,s_{m}^{\ell},s_{1}^{u},\ldots,s_{m}^{u})^{\prime},$
where $\underline{k}_{i}\equiv\operatorname{argmin}_{k\in\{1,\ldots,J_{U}\}}\{\tilde{u}_{k}+u_{k}(\hat{p}+\varepsilon_{i})\}$, $\bar{k}_{i}\equiv\operatorname*{argmax\;}_{k\in\{1,\ldots,J_{L}\}}\{\tilde{\ell}_{k}+\ell_{k}(\hat{p}+\varepsilon_{i})\}$, $s_{i}^{\ell}\equiv\operatorname*{sign}(\tilde{\ell}_{\bar{k}_{i}}+\ell_{\bar{k}_{i}}(\hat{p}+\varepsilon_{i}))$ and $s_{i}^{u}\equiv\operatorname*{sign}(\tilde{u}_{\underline{k}_{i}}+u_{\underline{k}_{i}}(\hat{p}+\varepsilon_{i}))$ for $i=1,\ldots,m$.
5 Confidence Interval Construction
We now generalize the CI construction described in Section 2 to apply in the general framework of Section 3, covering all example applications discussed above and in Appendix A. We start with conditional CIs and then move to unconditional CIs.
5.1 Conditional Confidence Intervals
We first generalize the conditional CI construction described in Section 2.2. As in Section 2.2, we are interested in forming probabilistic lower and upper bounds $\widehat{L}(\hat{d})_{\alpha}^{C}$ and $\widehat{U}(\hat{d})_{1-\alpha}^{C}$ that satisfy (2.2) and (2.5) for all $d\in\{d^{0},\ldots,d^{K}\}$ as endpoints in the formation of a conditionally valid CI. This is because the researcher’s interest in inference on option $d$ only arises when it is a member of the estimated set $\widehat{\mathcal{D}}$.
To begin, we characterize the conditional distributions of $\widehat{L}(d)$ given the event $\{d\in\widehat{\mathcal{D}}$, $\hat{j}_{L}(d)=j_{L}^{*}$ and $\hat{\gamma}_{L}(d)=\gamma_{L}^{*}\}$ characterized by Assumption 3.2. These conditional distributions depend upon the nuisance parameter $p$. As a first step, we form sufficient statistics for $p$ that are asymptotically independent of $\widehat{L}(d)$ given $\hat{j}_{L}(d)=j_{L}^{*}$ and $\widehat{U}(d)$ given $\hat{j}_{U}(d)=j_{U}^{*}$. Since $\widehat{L}(d)=\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p}$ given $\hat{j}_{L}(d)=j_{L}^{*}$ by Assumption 3.1 and (3.1) (and similarly for $\widehat{U}(d)$), such sufficient statistics can be constructed as
$\widehat{\mathcal{Z}}_{L}(d,j_{L}^{*})\equiv\sqrt{n}\hat{p}-\hat{b}_{L}(d,j_{L}^{*})\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p})$
@@ 5.1 Conditional Confidence Intervals
$\widehat{\mathcal{Z}}_{U}(d,j_{U}^{*})\equiv\sqrt{n}\hat{p}-\hat{b}_{U}(d,j_{U}^{*})\sqrt{n}(\tilde{u}_{d,j_{U}^{*}}+u_{d,j_{U}^{*}}\hat{p})$
@@
$\hat{b}_{L}(d,j_{L}^{*})\equiv\widehat{\Sigma}\ell_{d,j_{L}^{*}}^{\prime}\left(\ell_{d,j_{L}^{*}}\widehat{\Sigma}\ell_{d,j_{L}^{*}}^{\prime}\right)^{-1}\quad\text{and}\quad\hat{b}_{U}(d,j_{U}^{*})\equiv\widehat{\Sigma}u_{d,j_{U}^{*}}^{\prime}\left(u_{d,j_{U}^{*}}\widehat{\Sigma}u_{d,j_{U}^{*}}^{\prime}\right)^{-1},$
by the asymptotic normality of $\hat{p}$ and asymptotic absence of correlation between $\sqrt{n}\hat{p}$ and $\widehat{\mathcal{Z}}_{L}(d,j_{L}^{*})$ and $\widehat{\mathcal{Z}}_{U}(d,j_{U}^{*})$. Next, Assumption 3.2 characterizes the conditioning events $\{d\in\widehat{\mathcal{D}},\hat{j}_{L}(d)=j_{L}^{*},\hat{\gamma}_{L}(d)=\gamma_{L}^{*}\}$ and $\{d\in\widehat{\mathcal{D}},\hat{j}_{U}(d)=j_{U}^{*},\hat{\gamma}_{U}(d)=\gamma_{U}^{*}\}$ as polyhedra in $\hat{p}$, which can in turn be expressed as intervals for ~$\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p}$ and $\tilde{u}_{d,j_{U}^{*}}+u_{d,j_{U}^{*}}\hat{p}$; see, e.g., (3.10) and (3.19). Then, again since $\widehat{L}(d)=\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p}$ given $\hat{j}_{L}(d)=j_{L}^{*}$ by Assumption 3.1, Lemma 1 of McCloskey, (2024) implies
@@
$\sim\left.\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p})\right|\left\{\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d,j_{L}^{*}),d,j_{L}^{*},\gamma_{L}^{*}\right)\leq\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p})\leq\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}(d,j_{L}^{*}),d,j_{L}^{*},\gamma_{L}^{*}\right),\right.$
$\left.\widehat{\mathcal{V}}_{L}^{0}\left(\widehat{\mathcal{Z}}_{L}(d,j_{L}^{*}),d,j_{L}^{*},\gamma_{L}^{*}\right)\geq 0\right\}$ (5.1)
@@
$\sim\left.\sqrt{n}(\tilde{u}_{d,j_{U}^{*}}+u_{d,j_{U}^{*}}\hat{p})\right|\left\{\widehat{\mathcal{V}}_{U}^{-}\left(\widehat{\mathcal{Z}}_{U}(d,j_{U}^{*}),d,j_{U}^{*},\gamma_{U}^{*}\right)\leq\sqrt{n}(\tilde{u}_{d,j_{U}^{*}}+u_{d,j_{U}^{*}}\hat{p})\leq\widehat{\mathcal{V}}_{U}^{+}\left(\widehat{\mathcal{Z}}_{U}(d,j_{U}^{*}),d,j_{U}^{*},\gamma_{U}^{*}\right),\right.$
$\left.\widehat{\mathcal{V}}_{U}^{0}\left(\widehat{\mathcal{Z}}_{U}(d,j_{U}^{*}),d,j_{U}^{*},\gamma_{U}^{*}\right)\geq 0\right\}$
@@
$\widehat{\mathcal{V}}_{M}^{-}(z,d,j,\gamma)\equiv\max_{k:(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}<0}\frac{\sqrt{n}(c^{M}(d,j,\gamma))_{k}-(A^{M}(d,j,\gamma)z)_{k}}{(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}},$
$\widehat{\mathcal{V}}_{M}^{+}(z,d,j,\gamma)\equiv\min_{k:(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}>0}\frac{\sqrt{n}(c^{M}(d,j,\gamma))_{k}-(A^{M}(d,j,\gamma)z)_{k}}{(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}},$
$\widehat{\mathcal{V}}_{M}^{0}(z,d,j,\gamma)\equiv\min_{k:(A^{M}(d,j,\gamma)\hat{b}_{M}(d,j))_{k}=0}\sqrt{n}(c^{M}(d,j,\gamma))_{k}-(A^{M}(d,j,\gamma)z)_{k}$
@@
Now, under Assumptions 3.4 and 3.5, the distribution of $\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}\hat{p})$ can be approximated by a $\mathcal{N}(\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}p),\ell_{d,j_{L}^{*}}\widehat{\Sigma}\ell_{d,j_{L}^{*}}^{\prime})$-distributed random variable that is asymptotically independent of $\widehat{\mathcal{Z}}_{L}(d,j_{L}^{*})$. Using the distributional characterization in (5.1), we can therefore use the corresponding truncated normal cumulative distribution function to produce quantile-unbiased estimators of the underlying mean $\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}p)$. Let $F_{TN}(\cdot;\mu,\sigma^{2}|\mathcal{V}^{-},\mathcal{V}^{+})$ denote the truncated normal cumulative distribution function for an underlying normally-distributed random variable with mean $\mu$ and variance $\sigma^{2}$ that is truncated to lie between $\mathcal{V}^{-}$ and $\mathcal{V}^{+}$. For $\alpha\in(0,1)$, define $\widehat{L}(d)_{\alpha}^{C}$ to solve
$F_{TN}\left(\sqrt{n}(\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p});\mu,\ell_{d,\hat{j}_{L}(d)}\widehat{\Sigma}\ell_{d,\hat{j}_{L}(d)}^{\prime}\right|\widehat{\mathcal{V}}_{L}^{-}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right),$
$\left.\widehat{\mathcal{V}}_{L}^{+}\left(\widehat{\mathcal{Z}}_{L}(d,\hat{j}_{L}(d)),d,\hat{j}_{L}(d),\hat{\gamma}_{L}(d)\right)\right)=1-\alpha$
@@
$F_{TN}\left(\sqrt{n}(\tilde{u}_{d,\hat{j}_{U}(d)}+u_{d,\hat{j}_{U}(d)}\hat{p});\mu,u_{d,\hat{j}_{U}(d)}\widehat{\Sigma}u_{d,\hat{j}_{U}(d)}^{\prime}\right|\widehat{\mathcal{V}}_{U}^{-}\left(\widehat{\mathcal{Z}}_{U}(d,\hat{j}_{U}(d)),d,\hat{j}_{U}(d),\hat{\gamma}_{U}(d)\right),$
$\left.\widehat{\mathcal{V}}_{U}^{+}\left(\widehat{\mathcal{Z}}_{U}(d,\hat{j}_{U}(d)),d,\hat{j}_{U}(d),\hat{\gamma}_{U}(d)\right)\right)=1-\alpha$
in $\mu$. Then, results in Pfanzagl, (1994) imply that $\widehat{L}(d)_{\alpha}^{C}$ and $\widehat{U}(d)_{\alpha}^{C}$ are optimal $\alpha$ quantile-unbiased estimators of $\sqrt{n}(\tilde{\ell}_{d,j_{L}^{*}}+\ell_{d,j_{L}^{*}}p)$ and $\sqrt{n}(\tilde{u}_{d,j_{L}^{*}}+u_{d,j_{L}^{*}}p)$ asymptotically.
@@
Suppose Assumptions 3.1, 3.2 and 3.43.6 hold. Then, for any $d\in\{d^{0},\ldots,d^{K}\}$ and $20<\alpha_{1},\alpha_{2}<1/$ 2,
$\liminf_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}_{n}}\left\{\left[\mathbb{P}\left(\left.[L(d),U(d)]\subseteq\left(n^{-1/2}\widehat{L}(d)_{\alpha_{1}}^{C},n^{-1/2}\widehat{U}(d)_{1-\alpha_{2}}^{C}\right)\right|d\in\widehat{\mathcal{D}}\right)-(1-\alpha_{1}-\alpha_{2})\right]\cdot\mathbb{P}(d\in\widehat{\mathcal{D}})\right\}\geq 0$
for all $d\in\{d^{0},\ldots,d^{K}\}$.
5.2 Unconditional Confidence Intervals
In parallel with the previous subsection, we now generalize the unconditional CI constructions described in Section 2.3 to the general framework of Section 3. Note that conditional inference on $W(d)$ is well-defined for any given $d\in\widehat{\mathcal{D}}$, when conditioning on $d\in\widehat{\mathcal{D}}$. In contrast, unconditional inference on a data-dependent $W(d)$ requires it to be uniquely defined, as $W(\hat{d})$ in our notation. This is implied by Assumption 3.3. Here, we would like to construct CIs that unconditionally cover the identified interval corresponding to a unique data-dependent object of inferential interest. As mentioned in Section 2.3, if only unconditional coverage of $[L(\hat{d}),U(\hat{d})]$ is desired the conditional CI (5.2) with $d=\hat{d}$ can be unnecessarily wide. We describe two different methods—projection and hybrid methods—to form the unconditional probabilistic bounds that constitute the endpoints of these unconditional CIs in this general framework.
The general formation of the probabilistic bounds based upon projecting simultaneous confidence bounds for all possible values of $\sqrt{n}L(\hat{d})$ and $\sqrt{n}U(\hat{d})$ proceeds by computing $\hat{c}_{1-\alpha,M}$, the $1-\alpha$ quantile of $\max_{i\in\{1,\ldots,(T+1)J_{M}\}}\hat{\zeta}_{M,i}/\sqrt{\widehat{\Sigma}_{M,i}}$, where $\hat{\zeta}_{M}\sim\mathcal{N}(0,\widehat{\Sigma}_{M})$ for $M=L,U$ with $\widehat{\Sigma}_{L}=\ell^{mat}\widehat{\Sigma}\ell^{mat\prime}$, $\widehat{\Sigma}_{U}=u^{mat}\widehat{\Sigma}u^{mat\prime}$, $\ell^{mat}=(\ell_{0,1}^{\prime},\ldots,\ell_{0,J_{L}}^{\prime},\ldots,\ell_{T,1}^{\prime},\ldots,\ell_{T,J_{L}}^{\prime})^{\prime}$ and $u^{mat}=(u_{0,1}^{\prime},\ldots,u_{0,J_{U}}^{\prime},\ldots,u_{T,1}^{\prime},\ldots,u_{T,J_{U}}^{\prime})^{\prime}$, recalling that $\Upsilon_{i}$ denotes the $i^{th}$ element of the main diagonal of any square matrix $\Upsilon$. Here, the maximum is taken to guarantee simultaneous coverage. The lower level $1-\alpha$ projection confidence bound for $\sqrt{n}L(\hat{d})$ is $\widehat{L}(\hat{d})_{\alpha}^{P}=\sqrt{n}\widehat{L}(\hat{d})-\hat{c}_{1-\alpha,L}\sqrt{\widehat{\Sigma}_{L,\hat{d}J_{L}+\hat{j}_{L}(\hat{d})}}$ and the upper level $1-\alpha$ projection confidence bound for $\sqrt{n}U(\hat{d})$ is $\widehat{U}_{1-\alpha}^{P}(\hat{d})=\sqrt{n}\widehat{U}(\hat{d})+\hat{c}_{1-\alpha,U}\sqrt{\widehat{\Sigma}_{U,\hat{d}J_{U}+\hat{j}_{U}(\hat{d})}}$, because e.g., $\sqrt{n}L(\hat{d})$ can take value equal to any entry of the vector $\sqrt{n}(\tilde{\ell}_{0,1},\ldots,\tilde{\ell}_{0,J_{L}},\ldots,\tilde{\ell}_{T,1},\ldots,\tilde{\ell}_{T,J_{L}})^{\prime}+\sqrt{n}\ell^{mat}p$, $\sqrt{n}(\ell^{mat}\hat{p}-\ell^{mat}p)$ is asymptotically distributed $\mathcal{N}(0,\Sigma_{L})$ for $\Sigma_{L}=\ell^{mat}\Sigma\ell^{mat\prime}$ by Assumption 3.4 and $\widehat{\Sigma}_{L}$ is consistent for $\Sigma_{L}$ by Assumption 3.5.